Austin has 42 paperback books that are either Mysteries or Science Fiction. If he has 16 more mysteries than science fiction books, how many mysteries does he have?
Answer by CharStar(110)

Science Fiction Books = x
Mystery Book = x+16
Equation = x+x+16=42
2x=26
x=13
Science Fiction Books = 13
Mystery Books = 13+16= 29
Total = 42
I hope that helps you
CharStar